# Complementarity in Human-AI Collaboration: Concept, Sources, and Evidence

### v1

@@ 3 Conceptualization of Human-AI Complementarity
In this section, we first introduce the fundamental notion of human-AI complementarity, and formalize our decision-making situation as a basis for further analysis. Subsequently, we introduce and formalize the complementarity potential concept and discuss its possible sources.
@@ 3.1 The Principle of Human-AI Complementarity
We first motivate the underlying idea of complementarity which drives effective human-AI collaboration. In this work, we focus on the performance on decision-making tasks that humans and AI can conduct independently. We therefore recognize that AI contributions could, in many domains, go beyond mere (partial) decision support for humans, and could be regarded as an alternative way to solve a task independently. Real-world examples include diagnosing diseases in medicine (Goldenberg et al.,, 2019), conducting loan decisions in finance (Turiel and Aste,, 2020), and writing entire programs with AI code generation systems (Ross et al.,, 2023). However, since neither humans nor AI are perfect, their different available information or capabilities could be combined to generate superior outcomes in a human-AI team. Figure 1 illustrates the situation in a simple example for a set of binary decisions: The decision-making task comprises 25 instances with binary “right” or “wrong” outcomes. The number of incorrect decisions measures the performance. The AI makes 13 incorrect decisions, while the human errs at 15 task instances when conducting the task independently. Of all the instances, neither the AI nor the human can solve 5 on their own. Consequently, if the “human-AI team” were, for each instance, to pick the correct decision of either the AI or the human, the team could correctly solve 20 instances and would only miss the 5 that none of them can solve. In other words, human-AI collaboration opens an inherent “complementarity potential” of 8 instances compared to the individually better performing team member (here, the AI with 13 misses).
Figure 1: Illustration of the principle of human-AI complementarity.
@@
In reality, and depending on the application context, performance could be captured by more intricate measures than just the absolute number of errors, e.g., as the precision or the recall in classification tasks (e.g., when analyzing radiology images in health care), as the mean absolute error in prediction tasks (e.g., when making sales forecasts for inventory management), or as more complex compound metrics (e.g., when weighting multiple dimensions of interest). For complementarity potential to exist, it is essential that humans and AI have different strengths and weaknesses, i.e., that they make different errors across a variety of tasks (Geirhos et al.,, 2021; Steyvers et al.,, 2022). If their capabilities could be combined adequately, the team performance in such situations would be superior to their individual performances (Rastogi et al.,, 2023). In the following, we define this potential and its components formally.
@@ 3.2 Human-AI Decision-Making Setting
Let us first define the human-AI decision-making setting illustrated in Figure 1, which is the foundation of this work: A decision task $T=\left\{(x^{(i)},\ y^{(i)})\right\}_{i}^{N}$ is a set of $N$ instances $x^{(i)}\in X$ with corresponding ground truth labels $y^{(i)}\in Y$ denoting the correct results. Each instance $x^{(i)}$ represents an individual decision. In this section, we use the term decision for both classification and prediction tasks. The ground truth, i.e., the correct decision, might not be known at the time of the decision, but can be determined and revealed later. Both a human decision-maker $H$ and a machine learning model, which we denote as $AI$, are capable of independently producing a decision for each task instance. For any given instance $x^{(i)}$, the human and the AI will independently derive decisions $\hat{y}_{H}^{(i)}$ and $\hat{y}_{AI}^{(i)}$. In a scenario where the human and the AI might collaborate in respect of their decision-making, a collaboration mechanism $I(\hat{y}_{H}^{(i)},\hat{y}_{AI}^{(i)})$ combines their decisions into a final team decision $\hat{y}_{I}^{(i)}$. We note that this decision might be different from each individual decision. Each decision’s quality is measured by its deviation (“loss”) from the ground truth—by a loss function $l$ bounded in $R^{+}$. This function serves as a generic measure of performance that could take different forms with different decision problems, e.g., an error rate used in classification tasks (like the number of wrong or unsolved task instances in Figure 1). For any instance $x^{(i)}$, losses are given by $l_{H}^{(i)}$ for the human decision, $l_{AI}^{(i)}$ for the AI decision, and $l_{I}^{(i)}$ for the combined team decision. This results in overall losses $L_{D}$ for the entire task by averaging all the available instances:
$L_{D}=\frac{1}{N}\sum_{i=1}^{N}{l_{D}^{(i)}\left({\hat{y}}_{D}^{(i)},y^{(i)}\right)}\ with\ D\ \in\ \left\{H,AI,I\right\}$ (1)
@@ 3.3 Complementarity Potential
From a decision-theoretic perspective, the vision of human-AI collaboration is to attain a superior team performance compared to the human and the AI conducting the task individually—providing the fundamental reason for forming human-AI teams (Rastogi et al.,, 2023). In our context, the human-AI team reaches complementary team performance ($CTP$) when the loss of the team is strictly smaller than that of the human and the AI individually (Bansal et al.,, 2021):
@@
In addition to a binary task outcome, we propose the notion of complementarity potential ($CP$) to measure the discrepancy between the overall loss of the individually better performing team member $T^{\ast}\in\{H,AI\}$ with $L_{T^{\ast}}=\min\left(L_{H},L_{AI}\right)$ and perfect decisions for all instances of a task, i.e., selecting the ground truth, with an overall loss of 0:
$CP=\ L_{T^{\ast}}=min\left(L_{H},L_{AI}\right)$ (3)
@@
In our introductory example in Figure 1, the overall loss is quantified by the number of wrong decisions. Consequently, the complementarity potential amounts to 13, which is given as the minimum number of individual errors (13 for the AI, 15 for the human). [^2] For simplification, we report $L_{D}$ in the introductory example as the sum instead of the average of all the instances. It should be noted that only 8 task instances of this potential can be realized by picking the better individual decision, while 5 task instances cannot be solved individually, but only—if at all—in collaboration. We reflect this by distinguishing two $CP$ components: inherent and collaborative complementarity potential.
@@
Inherent complementarity potential represents improvements, i.e., loss reductions, that—from the perspective of the overall more accurate team member $T^{\ast}$—could be contributed by including any superior decisions on the instance level by the overall less accurate team member. Figure 2 illustrates this—assuming, without loss of generality, that AI is the overall better performing individual team member ($L_{AI}\leq L_{H}$). In respect of task instances where the inferior team member, i.e., the human, could help to reduce the loss, we capture the inherent complementarity potential.
@@
Figure 2: Complementarity potential ($CP$) split into inherent ($CP^{inh}$) and collaborative ($CP^{coll}$) complementarity potential for a single instance with better human performance (left) and better AI performance (right) in respect of $T^{\ast}=AI$. $l_{D}$ denotes the instance-specific loss with $D\in\ \left\{H,AI\right\}$.
The remaining losses constitute the collaborative complementarity potential that could only be exploited if the team members’ collaboration yields new insights not available for the individual decisions before the collaboration.
@@
Formally, inherent ${CP}^{inh}$ can be calculated by aggregating all the potential loss improvements for the overall better performing team member across all instances:
${CP}^{inh}=\begin{cases}\frac{1}{N}\sum\limits_{i=1}^{N}\max(0,l_{AI}^{(i)}-l_{H}^{(i)}),&L_{AI}\leq L_{H},\\ \frac{1}{N}\sum\limits_{i=1}^{N}\max(0,l_{H}^{(i)}-l_{AI}^{(i)}),&L_{AI}>L_{H}.\end{cases}$ (4)
@@
The remaining collaborative ${CP}^{coll}$ can be calculated by aggregating the remaining minimum losses that the team members incur individually per task instance:
${CP}^{coll}=\frac{1}{N}\sum_{i=1}^{N}\min(l_{H}^{(i)},l_{AI}^{(i)})$ (5)
Collaborative potential signifies an improvement potential that goes beyond individual solutions in generating “new” knowledge. In respect of classification tasks, this could entail identifying the correct class decisions that neither match the (incorrect) human or AI decisions—based on the disagree of both. In respect of regression tasks, over- and underestimates of the respective team members may result in averaged estimates that are more accurate than the individual ones.
@@
The inherent and collaborative components are additive and together form the complementarity potential (see Appendix A for additional details):
$CP={CP}^{inh}+{CP}^{coll}$ (6)
@@
In our introductory example, the total complementarity potential of 13 instances can be differentiated into inherent complementarity potential (${CP}^{inh}$) of 8 instances (for which the human team member can contribut the correct solution), and a collaborative complementarity potential (${CP}^{coll}$) of 5, with none of the team members arriving at the correct decision individually.
Figure 3: Illustration of (theoretical) complementarity potential and (realized) complimentarity effect extending the introductory example.
3.4 Realization of Complementarity Potential
@@
In real world collaboration scenarios between a human and AI, it is, of course unlikely that the entire complementarity potential will be exploited. We therefore introduce the complementarity effect ($CE$) as the part of this potential that is realized by the joint team decision. Measuring and dissecting this effect will allow observed human-AI team settings to be analyzed in greater detail, in order to infer conclusions about the collaboration’s effectiveness, and to purposefully develop and compare collaboration mechanisms. Analogous to the complementarity potential in Equation 3, the realized complementarity effect denotes the difference between the average loss of the overall individually better team member and that of the joint team decision ($L_{I}>0$):
$CE\ =\ min(L_{H},L_{AI})-L_{I}$ (7)
@@
Figure 3 extends our introductory example from Figure 1 by additionally incorporating hypothetically realized human-AI decisions for all instances. We find that the human-AI team only makes 9 incorrect decisions compared to the 13 of the AI and the 15 of the human, when they act independently. This collaboration therefore reaches a complementarity effect of 4—realizing 31% of the full complementarity potential of 13.
Figure 4: Illustration of (theoretical) complementarity potential ($CP$) and (realized) complimentarity effect ($CE$) for different loss scenarios on an instance level—assuming, without loss of generality, that the AI performs better overall ($T^{\ast}=AI$). $l_{D}$ denotes the instance-specific loss with $D\in\ \left\{H,AI,I\right\}$.
@@
The complementarity effect measures loss improvements—from the perspective of the overall more accurate team member $T^{\ast}$—that are realized by a particular combination of the individual decisions into a team decision. Analogous to inherent or collaborative potential, we can split the complementarity effect into the same two categories. Figure 4 illustrates this—again, we assume, without loss of generality, the AI to be the overall better performing individual team member ($L_{AI}\leq L_{H}$). If there is inherent potential (i.e., for task instances where the overall inferior human is more knowledgeable, as in scenario 1-3) this inherent complementarity potential may be realized partially ($l_{AI}>l_{I}>l_{H}$, scenario 1), fully ($l_{AI}>l_{H}>l_{I}$, scenario 2), or not at all ($l_{I}>l_{AI}$, scenario 3). In scenario 2, any improvement beyond the individual loss of the team member who is more accurate for a particular instance means that also the collaborative effect occurs. If there is no inherent complementarity effect (as in scenarios 3 and 4), any realization effect counts to the collaborative effect. We note, however, that the collaborative effect could also be negative, e.g., for $l_{I}>l_{AI}$ as in scenario 3. This would indicate that in these scenarios the collaboration rather deteriorates the outcomes. [^3] Note, that the negative collaborative effect should not be interpreted as unrealized inherent complementarity potential, as it occurs independently of the overall inferior team member’s individual loss.
In general, we can aggregate the complementarity effects across all instances of a task and summarize both cases (AI or human with overall better performance) and the scenarios above (depending on the instance performance of human, AI, and human-AI team):
${CE}^{inh}$ $=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}l_{AI}^{(i)}-l_{I}^{(i)},&L_{AI}\leq L_{H}\text{ and }l_{AI}^{(i)}>l_{I}^{(i)}\geq l_{H}^{(i)},\\ l_{AI}^{(i)}-l_{H}^{(i)},&L_{AI}\leq L_{H}\text{ and }l_{AI}^{(i)}>l_{H}^{(i)}>l_{I}^{(i)},\\ l_{H}^{(i)}-l_{I}^{(i)},&L_{H}<L_{AI}\text{ and }l_{H}^{(i)}>l_{I}^{(i)}\geq l_{AI}^{(i)},\\ l_{H}^{(i)}-l_{AI}^{(i)},&L_{H}<L_{AI}\text{ and }l_{H}^{(i)}>l_{AI}^{(i)}>l_{I}^{(i)},\\ 0,&\text{otherwise}.\end{cases}$ (8)
${CE}^{coll}$ $=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}l_{H}^{(i)}-l_{I}^{(i)},&l_{AI}^{(i)}\geq l_{H}^{(i)}>l_{I}^{(i)},\\ l_{AI}^{(i)}-l_{I}^{(i)},&l_{H}^{(i)}\geq l_{AI}^{(i)}>l_{I}^{(i)},\\ l_{AI}^{(i)}-l_{I}^{(i)},&L_{AI}\leq L_{H}\text{ and }l_{I}^{(i)}>l_{AI}^{(i)},\\ l_{H}^{(i)}-l_{I}^{(i)},&L_{H}<L_{AI}\text{ and }l_{I}^{(i)}>l_{H}^{(i)},\\ 0,&\text{otherwise}.\end{cases}$ (9)
@@
Figure 5 summarizes the complementarity potential, including its realization. It allows us to assess the collaboration between humans and AI more comprehensively. We will later illustrate the application of this concept in our experiments. Beforehand we, however, contemplate on the sources underlying the potential for complimentarity.
Figure 5: Overview of the complementarity potential and effect, including the inherent and collaborative components. Note that performance is represented by a loss that can be interpreted as an error measure. This means a lower value represents a better performance.
3.5 Sources of Complementarity Potential
A continuously growing body of research on human-AI decision-making assumes there are complementary capabilities between humans and AI (Bansal et al.,, 2021; Dellermann et al.,, 2019). The discourse is, however, held on a general level, e.g., by hypothesizing that humans excel at creativity, whereas AI better identifies patterns in large amounts of data (Dellermann et al.,, 2019). In this context, a conceptualization and empirical analysis of complementarity are still lacking. We therefore aim to provide a more nuanced understanding of complementarity potential by identifying and structuring its sources. To this end, we conceptually distinguish three phases in human-AI decision-making, as shown in Figure 6.
The first is the learning or training phase encompassing AI training and the human learning process. Note that AI models are usually developed within a short period of time through learning from aggregated training data, while human learning is a lifelong process. The second is the inference phase, in which the human and AI infer a decision pertaining to a particular instance. The third is the collaboration phase, in which the human and AI interact. Based on this simplified decision-making process, we discuss the sources of complementarity potential more granularly. We hypothesize that complementarity potential is created through asymmetries in either the information or the capabilities.
Training Phase.
First, training data as input of the training process differs between humans and AI. Humans might have experienced a considerable number of training instances, whereas AI is usually trained on a limited data set. Second, humans and AI’s inherent capabilities, which training stimulates, differ from one another. AI is, for instance, able to efficiently identify patterns in high-dimensional data or infer decisions from probabilistic reasoning (Dellermann et al.,, 2019; Jarrahi,, 2018). In contrast, humans are already able to learn abstract concepts from a small number of samples (Zheng et al.,, 2017). In this context, human decision-making is often purely heuristic or intuitive and does not consider all of the available information (Jarrahi,, 2018). Based on these asymmetries during the training process, humans and AI learn different decision boundaries from which their individual decisions are inferred (Geirhos et al.,, 2020, 2021; Steyvers et al.,, 2022). These differences could materialize in distinct capabilities; their presence results in inherent complementarity potential.
Figure 6: Conceptual overview of the human-AI decision-making process.
Inference Phase.
After the training process, the AI can infer a decision for a particular instance. Even when assuming that humans and AI have identical decision boundaries, different available information can constitute a source of inherent complementarity potential—in real-world settings, AI and humans often have access to different features (Bansal et al.,, 2019; Sanders and Ritzman,, 2001). A famous example is the “broken leg” scenario, which refers to side information known to humans (Meehl,, 1957). This information could not be incorporated as features into the model due to its rare occurrence. In many application domains in which AI is applied to support human decision-making, there is additional information beyond the data used to train the AI model (Ibrahim et al.,, 2021). In practice, due to technical or economic reasons, this discretionary data may often not be digitally available at all, or only in a small quantity insufficient for model training (Ibrahim et al.,, 2021). Nevertheless, in human-AI decision-making, human team members might leverage their expertise to use the additional information (Ibrahim et al.,, 2021). On the other hand, AI models could also have access to unique contextual information that is unavailable to human team members during the decision-making process.
To summarize, inherent complementarity potential can be present during training or inference and is usually either based on information or capabilities asymmetry.
Collaboration Phase.
The collaboration process adds a third phase to the overall decision-making process. In this phase, collaborative complementarity potential might arise as an additional part of complementarity potential beyond the inherent complementarity potential. For the collaborative aspect of complementarity potential, it is essential for the human-AI team to possess the capability to further improve the best individual decision achieved by either the AI or human in respect of a given task instance.
In the following, we illustrate the effects of information and capability asymmetry in behavioral experiments by applying the complimentarity concepts developed in this section.

### v2

@@ 3 Conceptualization of Human-AI Complementarity
In this section, we first introduce the fundamental notion of human-AI complementarity and formalize our decision-making situation as a basis for further analysis. Subsequently, we introduce and formalize the concepts of complementarity potential and effect, and relate them to the underlying sources of complementarity.
@@ 3.1 The Principle of Human-AI Complementarity
We first motivate the underlying idea of complementarity which drives effective human-AI decision-making. In this work, we focus on the performance on decision-making tasks that humans and AI can conduct independently—recognizing that AI models have elevated above pure decision support for humans. Real-world examples include diagnosing diseases in medicine (Goldenberg et al., 2019), conducting loan decisions in finance (Turiel and Aste, 2020), and writing entire programs with AI code generation systems (Ross et al., 2023). However, since neither humans nor AI are perfect, discrepancies in access to information or in capabilities could be leveraged to generate superior outcomes in a human-AI team. Figure 2 illustrates the situation in a simple example for a set of decisions: The decision-making task comprises 25 instances that each have a set of possible discrete outcomes only one of which is correct. The number of incorrect decisions measures the performance of each individual team member and the human-AI team, respectively. The AI makes 13 incorrect decisions, while the human errs at 15 task instances when conducting the task independently. Five of the instances can neither be solved by the AI nor the human on their own. Consequently, if the human-AI team were just to pick the correct decision of either the AI or the human, the team could correctly solve 20 instances and would only miss the 5 that none of them can solve. In other words, relying on the correct individual decisions of each team member improves the result compared to the AI acting alone (as the individually better performing team member): While the AI still shows 13 incorrect decisions when acting independently, teaming up with the human can reduce this number to 5—realizing an improvement potential of 8 decisions. Moreover, it is also conceivable that—while none of the team members is correct individually—the interaction between the team members may allow the generation of correct team decisions even for the remaining 5 red task instances in Figure 2 representing an additional improvement potential [^2] We note that this is not possible if team members rely on one of their individual decisions as team decision—as tasks solutions are confined to the solutions provided by each team member alone.: Team members may recognize through collaboration that they have different information or capabilities that they can jointly apply. Let us, e.g., assume that a task would be to deliver a diagnosis on cancer. Both a human radiologist and an AI analyzing patient data might each individually render a wrong diagnosis, e.g., “malign cancer” conjectured by the human and “no cancer” proposed by the AI. If, however, the human might learn about the AI’s rationale (e.g., regions in the X-ray scan crucial for the decision) or the AI might learn about “side” information that the radiologist has on the patient history, they jointly may arrive at the correct diagnosis “benign cancer”.
For human-AI teams to achieve CTP, it is essential that they manage to realize these improvement potentials. If their information and capabilities could be adequately combined, the team performance in such situations would be superior to their individual performances (Rastogi et al., 2023).
@@
In the introductory example, performance is captured by the absolute number of errors. Depending on the application context, more intricate measures could also be applied, e.g., precision or recall in classification tasks (e.g., when analyzing radiology images in health care), mean absolute error in prediction tasks (e.g., when making sales forecasts for inventory management), or more complex compound metrics (e.g., when weighting multiple dimensions of interest).
Figure 2: Illustration of the principle of human-AI complementarity based on different (in)correct decisions that human and AI can make.
In reality, the performance of humans and AI varies depending on the task and the application domain. Enabled by advances in AI research over the last few years, there has been an increasing number of tasks where the performance of AI has reached or even exceeded that of humans (Afshar et al., 2022; Bubeck et al., 2023; Silver et al., 2018). However, many applications remain in which human performance remains superior (Brynjolfsson et al., 2018; Grace et al., 2018, 2024). For the motivating example, we chose a setting in which the AI makes fewer errors—as a reflection of the developments in AI model performance over the last years. However, the conceptualization that we develop in this section is independent of the performance relationship between human and AI.
@@ 3.2 Human-AI Decision-Making Setting
Let us first define the human-AI decision-making setting illustrated in Figure 2, which is the foundation of this work: A decision task $T=\left\{(x^{(i)},\ y^{(i)})\right\}_{i}^{N}$ is a set of $N$ instances $x^{(i)}\in X$ with corresponding ground truth labels $y^{(i)}\in Y$ denoting the correct results. In this section, we use the term decision for both classification and prediction tasks. The ground truth, i.e., the correct decision, might not be known at the time of the decision, but can be determined and revealed later. Both a human decision-maker H and a machine learning model, which we denote as AI, are capable of independently producing a decision for each task instance. For any given instance $x^{(i)}$, the human and the AI will independently derive decisions $\hat{y}_{H}^{(i)}$ and $\hat{y}_{AI}^{(i)}$. In a scenario where the human and the AI might collaborate in respect of their decision-making, a collaboration mechanism $I(\hat{y}_{H}^{(i)},\hat{y}_{AI}^{(i)})$ integrates their decisions into a final team decision $\hat{y}_{I}^{(i)}$. We note that this decision might be different from each individual decision.
Each decision’s quality is measured by its deviation (“loss”) from the ground truth—by a loss function $l$ bounded in $R^{+}$. This function serves as a generic measure of performance that could take different forms with different decision problems, e.g., an error rate used in classification tasks (like the number of wrong or unsolved task instances in Figure 2). For any instance $x^{(i)}$, losses $l_{D}^{(i)}$ with $D\in\left\{H,AI,I\right\}$ are given by $l_{H}^{(i)}$ for the human decision, $l_{AI}^{(i)}$ for the AI decision, and $l_{I}^{(i)}$ for the integrated team decision. This results in overall losses $L_{D}$ for the entire task by averaging all the available instances:
$L_{D}=\frac{1}{N}\sum_{i=1}^{N}{l_{D}^{(i)}\left({\hat{y}}_{D}^{(i)},y^{(i)}\right)}\ with\ D\ \in\ \left\{H,AI,I\right\}.$ (1)
@@ 3.3 Complementarity Potential
From a decision-theoretic perspective, the vision of human-AI decision making is to attain a superior team performance compared to the human and the AI conducting the task individually—providing the fundamental reason for forming human-AI teams (Rastogi et al., 2023). In our context, the human-AI team reaches complementary team performance (CTP) when the loss of the team is strictly smaller than that of the human and the AI individually (Bansal et al., 2021):
@@
In addition to a binary task outcome, we propose the notion of complementarity potential ($CP$) to measure the discrepancy between the overall loss of the individually better performing team member $T^{\ast}\in\{H,AI\}$ with $L_{T^{\ast}}=\min\left(L_{H},L_{AI}\right)$ and perfect decisions for all instances of a task, i.e., the selection of the ground truth, with an overall loss of 0:
$CP=\ L_{T^{\ast}}=min\left(L_{H},L_{AI}\right).$ (3)
@@
In our introductory example in Figure 2, the overall loss is quantified by the number of wrong decisions. Consequently, the complementarity potential amounts to 13, which is given as the minimum number of individual errors (13 for the AI, 15 for the human). [^3] For simplification, we report $L_{D}$ in the introductory example as the sum instead of the average of all the instances. It should be noted that only 8 task instances of this potential can be realized by picking the better individual decision, while 5 task instances cannot be solved individually, but only—if at all—in collaboration, resulting in a decision that neither AI nor human could come up with on their own. We reflect this by distinguishing two components: Inherent and collaborative complementarity potential.
@@
Inherent complementarity potential represents improvement potential, i.e., loss reductions, that—from the perspective of the overall more accurate team member $T^{\ast}$—could be contributed by any superior decisions on the instance level by the overall less accurate team member. This means that the team decision is confined to the solutions contributed by one of the team members—in the introductory example (Figure 2) represented by the 20 “gray and black” task instances solvable by at least human or AI:
$\hat{y}_{I}^{(i)}=I(\hat{y}_{H}^{(i)},\hat{y}_{AI}^{(i)})\in\left\{\hat{y}_{H}^{(i)},\hat{y}_{AI}^{(i)}\right\}.$ (4)
Collaborative complementarity potential, on the other hand, signifies improvement potential, i.e., loss reductions, that go beyond the individual team solutions by generating “new” knowledge. It may be noted that this is only possible if there is interaction between the team members enabling them to learn from the result of the partner and realize integrated values $\hat{y}_{I}^{(i)}\in Y$ different from the individual ones ($\hat{y}_{H}^{(i)}$ and $\hat{y}_{AI}^{(i)}$). In the introductory example (Figure 2), this is captured by the 5 “red” task instances not solvable by either team member alone.
We illustrate both potentials from a general perspective on a continuous range by looking at the individual losses of human and AI for a single task instance in Figure 3. Without loss of generality, we assume that the AI is the overall better performing individual team member ($L_{AI}\leq L_{H}$). If for this particular instance the overall inferior team member, i.e., the human, could help reduce the loss with his/her decision, we denote inherent complementarity potential (scenario 1). The remaining loss is unavoidable if—according to Equation 4—the team decision is restricted to one of the team members’ individual decisions. Thus, the loss of the better performing team member for a task instance constitutes the collaborative complementarity potential. This potential could only be exploited if the team members’ collaboration yields new insights for this task instance that were not available for the individual decisions before the collaboration (scenario 1 and 2), resulting in a team decision that incurs a loss lower than that of each team member (human and AI) individually.
@@
Figure 3: Complementarity potential ($CP$) split into inherent ($CP^{inh}$) and collaborative ($CP^{coll}$) components for a single instance with better human performance (left) and better AI performance (right) in respect of $T^{\ast}=AI$. $l_{D}$ denotes the instance-specific loss with $D\in\ \left\{H,AI\right\}$ with a lower loss indicating better performance for the same instance.
@@
Formally, inherent complementarity potential ${CP}^{inh}$ can be calculated by aggregating all potential loss improvements on the instance level where the overall worse performing team member can help improve the team result:
${CP}^{inh}=\frac{1}{N}\sum\limits_{i=1}^{N}\begin{cases}\max(0,l_{AI}^{(i)}-l_{H}^{(i)}),&L_{AI}\leq L_{H},\\ \max(0,l_{H}^{(i)}-l_{AI}^{(i)}),&L_{AI}>L_{H}.\end{cases}$ (5)
@@
The collaborative complementarity potential ${CP}^{coll}$ can be calculated by aggregating the remaining minimum losses per task instance that the team members incur individually:
${CP}^{coll}=\frac{1}{N}\sum_{i=1}^{N}\min(l_{H}^{(i)},l_{AI}^{(i)}).$ (6)
@@
The inherent and collaborative components are additive and together form the total complementarity potential $CP$ (see Appendix A for additional details):
$CP={CP}^{inh}+{CP}^{coll}.$ (7)
@@
In our introductory example, the total complementarity potential of 13 instances can be differentiated into an inherent complementarity potential (${CP}^{inh}$) of 8 instances (for which the human team member can contribute the correct solution), and a collaborative complementarity potential (${CP}^{coll}$) of 5, with none of the team members arriving at the correct decision individually.
3.4 Complementarity Effect
@@
In real-world collaboration scenarios between a human and AI, it is, of course, unlikely that the entire complementarity potential will be exploited. We, therefore, introduce the complementarity effect ($CE$) as that part of this potential that is actually realized by the integrated team decision. Measuring and dissecting this effect will allow observed human-AI decision-making settings to be analyzed in greater detail, in order to infer conclusions about the collaboration’s effectiveness, and to purposefully develop and compare collaboration designs and mechanisms. Analogous to the complementarity potential in Equation 3, the realized complementarity effect accounts for the difference between the average loss of the overall individually better team member and that of the integrated team decision ($L_{I}>0$):
$CE\ =\ min(L_{H},L_{AI})-L_{I}.$ (8)
@@
We expand our introductory example (Figure 2) in Figure 4 by incorporating (hypothetical) integrated human-AI decisions for all instances. Let us assume that the human-AI team makes 9 incorrect decisions, compared to 13 errors by the AI and 15 errors by the human when acting independently. Thus, this collaboration generates a complementarity effect of 4—realizing 31% of the full complementarity potential of 13.
Figure 4: Illustration of (theoretical) complementarity potential and (realized) complementarity effect for a hypothetical situation extending the introductory example in Figure 2.
@@
The complementarity effect measures loss improvements—from the perspective of the overall more accurate team member $T^{\ast}$—that are realized by the integration of individual decisions into a team decision. Analogous to inherent or collaborative complementarity potential, we can split the complementarity effect into the same two categories. Figure 5 illustrates the different scenarios that could materialize for a particular task instance in terms of the losses caused by the solutions of human, AI, and the (integrated) team for the same task instance. Again, we assume, without loss of generality, the AI to be the overall better performing individual team member ($L_{AI}\leq L_{H}$). If there is inherent complementarity potential (i.e., the overall inferior human is more knowledgeable for the particular task instance), this potential may be realized either partially ($l_{AI}>l_{I}>l_{H}$, scenario 1), fully ($l_{AI}>l_{H}>l_{I}$, scenario 2), or not at all ($l_{I}>l_{AI}$, scenario 3). No inherent complementarity potential exists where the overall better performing team member also dominates in the particular task instance (scenario 4).
Figure 5: Illustration of (theoretical) complementarity potential ($CP$) and (realized) complimentarity effect ($CE$) in respect of different loss scenarios for a single instance—assuming, without loss of generality, that the AI performs better overall ($T^{\ast}=AI$). $l_{D}$ denotes the instance-specific loss with $D\in\ \left\{H,AI,I\right\}$—with a lower loss indicating better performance for the same instance.
In addition, the collaborative complementarity potential (as the smaller loss of the team members in each scenario) could be tapped into. While in scenario 1 only a fraction of the inherent complementarity potential is realized, scenario 2 not only fully exploits the inherent complementarity potential, but also taps into some of the collaborative complementarity potential. Neither scenarios 3 nor scenario 4 realize any inherent complementarity potential, but solely contribute collaborative complementarity effects: In scenario 3 the integrated solution performs even worse than that of the inferior team member—resulting in a negative collaborative complementarity effect, i.e., the collaboration actually worsens the outcome. Conversely, in scenario 4, the integrated solution outperforms the better team member, generating a positive collaborative complementarity effect. In general, we can aggregate the complementarity effects across all instances of a task and summarize both cases (AI or human with overall better performance) and the scenarios above (depending on the instance performance of human, AI, and human-AI team):
${CE}^{inh}$ $=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}l_{AI}^{(i)}-l_{I}^{(i)},&L_{AI}\leq L_{H}\hskip 1.99997pt\text{ and }l_{AI}^{(i)}>l_{I}^{(i)}\geq l_{H}^{(i)},\\ l_{AI}^{(i)}-l_{H}^{(i)},&L_{AI}\leq L_{H}\hskip 1.99997pt\text{ and }l_{AI}^{(i)}>l_{H}^{(i)}>l_{I}^{(i)},\\ l_{H}^{(i)}\hskip 0.74997pt-l_{I}^{(i)},&L_{H}\hskip 2.25006pt<L_{AI}\text{ and }l_{H}^{(i)}>l_{I}^{(i)}\geq l_{AI}^{(in)},\\ l_{H}^{(i)}\hskip 0.74997pt-l_{AI}^{(i)},&L_{H}\hskip 2.25006pt<L_{AI}\text{ and }l_{H}^{(i)}>l_{AI}^{(i)}>l_{I}^{(i)},\\ 0,&\text{otherwise}.\end{cases}$ (9)
[…]
@@
3.5 Sources of Complementarity: The Impact on Complementarity Potential and Effect
In the motivating example, the human and the AI make erroneous decisions for different task instances allowing for possible performance synergies that may be attainable through collaboration. We used this example to introduce the measures of complementarity potential and effect.
Differences in human and AI decision-making can arise from information and capability asymmetries as introduced earlier (Section 2.1). These asymmetries are fundamental for the existence of complementarity. They cause different decisions to be taken by humans and AI on an instance level and, therefore, influence the complementarity potential and the realized complementarity effect including both components. Leveraging information asymmetry means making information available to the team: Advantages can be captured as inherent (when the team decision for a task instance corresponds to the individual decision of the team member with “better” information) or as collaborative complementarity potential/effect (when the information of both team members is joined resulting in new insights and a team decision different from the individual ones). Similarly, capability asymmetries may contribute to inherent (when the team decision corresponds to the individual decision of the team member that is more capable for the particular task instance) or collaborative complementarity potential/effect (when capabilities complement each other, e.g., human experience and AI computational power, resulting in a team decision that differs from the individual ones).
It may be noted, though, that realizing complementarity potential (and, thus, achieving CTP) involves managing a number of trade-offs: Information and capabilities are typically not distributed in a way that either the human or the AI dominates across all existing task instances (Geirhos et al., 2020, 2021; Kühl, Goutier, Baier, Wolff and Martin, 2022). There may be pieces of information that only the human or only the AI has access to: In our earlier radiology example, this may be contextual information about a patient that the physician has and a large variety of cases in the training set that only the AI is able to access. In addition, also information may be traded off against capabilities: Assuming the radiologist has dominating information, it may be outweighed by the capabilities of the AI to automatically digest and evaluate the information available to it in real time.
Figure 6: Conceptual framework for human-AI complementarity. It summarizes the complementarity potential and effect, including the inherent and collaborative components. Information and capability asymmetry can affect both the theoretically existing potential and the realized effect.
We summarize the notion of complementarity, its sources, and the measures of complementarity potential and effect in the conceptual framework in Figure 6. Each source can affect both the “inherent” and “collaborative” components of complementarity potential and effect. Our conceptualization presented in this section is intended to provide a deeper understanding of the phenomenon as well as to provide a concrete measurement construct for systematically harnessing complementarity.
To illustrate and evaluate this framework, we now design two behavioral experiments in which humans and AI make decisions on their own and jointly for a set of task instances. We investigate how the presence of information (Section 4.1) and capability (Section 4.2) asymmetry affect the team’s complementarity potential, the realized complementarity effect as well as the joint team performance.